The velocity of light in a liquid is 0.80 as fast as it is in air. What is the index of




refraction of the liquid?

1
Expert's answer
2022-03-02T10:15:04-0500

speed of light in the liquid "=0.80\times 3\times 10^8=240000000m/s"

refractive index=speed of light in vacuum"\/" speed of light in liquid.


"=\frac{(3\times 10^8}{240000000}=1.25"